72cm³ of a mixture of Nitrogen and Nitrogen oxide were passed over heated copper until all the nitrogen oxide had been reduced to nitrogen. The volume of the resulting gas was 60cm³, calculate the volume of nitrogen oxide in the original mixture.

1
Expert's answer
2022-01-28T06:31:03-0500

2NO+2Cu"=" 2CuO+ N2


Let V"=" Volume of Nitrogen

Volume of NO "=72-V"

Volume of "N_2=60-V"


NO:N2 "=2:1"


"72-V=2(60-V)"

"V=48"



Volume of NO "=72-48=24cm^3"